OpenStudy (anonymous):

Has the polynomial been factored correctly? x^5-16x^3+8x^2-128 x^3(x^2-16)+8(x^2-16) (x^2-16)(x^3+8) If not explain why please

OpenStudy (anonymous):

wrong by try 8x^2 =! -16x^2

OpenStudy (ranga):

(x^2 - 16) is x^2 - 4^2 that can be factored further. Use the identity: a^2 - b^2 = (a+b)(a-b) And x^3 + 8 = x^3 + 2^3 which can also be factored further. Use the identity: a^3 + b^3 = (a+b)(a^2 - ab + b^2)